OpenStudy (xo.msml):

Maria looks at the architectural plan of a four-walled room in which the walls meet each other at right angles. The length of one wall in the plan is 13 inches. The length of the diagonal of the floor of the room in the plan is approximately 15.26 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work. HELP!

OpenStudy (andrewishere):

No this is no a square. A square is perfectly equal in all lengths (width, length, height, diagonals, etc) This problem shows that the length of a wall is smaller than its diagonal, which is inconsistent with the description of a square, which means this shape is not a square.

OpenStudy (andrewishere):

Does this help?

OpenStudy (88):

if the floor is a square, then the other side will also be 13 right ?

OpenStudy (adi3):

yup to me

OpenStudy (anonymous):

you have 2 parts of a triangle for pythagorean theorem a^2+b+2=c^2 you have 13^2+b^2=15.26^2 solve for b^2

OpenStudy (88):